OpenStudy (anonymous):

Solve –3x2 – 4x – 4 = 0 need answers please using quadtratic equation

OpenStudy (anonymous):

do you know the quadratic equation? you're ready to go with ax2 + bx + c = 0

OpenStudy (anonymous):

i just need the answer

OpenStudy (anonymous):

type in "factor (-3x^2-4x-4 = 0)" into www.wolframalpha.com

OpenStudy (anonymous):

oops I mean "factor (-3x^2-4x-4)" or alternatively "(-3x^2-4x-4 = 0)" or even "solve -3x^2-4x-4 = 0" ... wolfram is pretty generous about your grammar

OpenStudy (anonymous):

answers are complex numbers (imaginary)

OpenStudy (mertsj):